# A Computationally Efficient Maximum A Posteriori Sequence Estimation via Stein Variational Inference

### v3

@@ IV-A Sequential Variational Inference for Dynamical Systems
Under Assumption IV.1, we can show that the optimal solution of (9) admits the following factorized structure.
@@
According to Lemma IV.1, the proposal distribution $q(x_{0:T})$ is factorized in (IV-A). Consequently, the ELBO in (IV-A) can be expressed as a summation of factorized KL divergence terms, each marginalized with respect to the previous optimal distribution, as stated in the following result.
@@ IV-B1 Viterbi-Style Dynamic Programming Algorithm
IV-B2 Theoretical Guarantees and Computational Complexity
@@
Moreover, it substantially reduces both computational and memory requirements, as detailed below.
Remark IV.1
(Computational and Space Complexities). The Viterbi algorithm requires $\mathcal{O}(N^{2})$ computational complexity and $\mathcal{O}(NT)$ space complexity for the memorization table, where $N$ is the number of discrete states. In general, achieving high accuracy necessitates a sufficiently large $N$. In contrast, Stein-MAP-Seq operates with a much smaller number of discrete states (particles), where $N_{s}\ll N$. Consequently, the computational complexity is reduced to $\mathcal{O}(N_{s}^{2})$ (with $N_{s}^{2}\ll N^{2}$) and the space complexity to $\mathcal{O}(N_{s}T)$ (with $N_{s}T\ll NT$). Furthermore, because the SVGD updates in (IV.3) can be computed in parallel, Stein-MAP-Seq is well-suited for efficient execution on parallel computing devices (e.g., GPUs or multi-core CPUs). This contrasts with Monte Carlo sampling methods such as particle filters [38], where the resampling step introduces synchronization bottlenecks. $\Box$
// 1) Initialization $(t=1)$
1 for $i=1,2,\dots,N_{s}$ in parallel do
2 $x_{1}^{i}\leftarrow x_{1}^{i}+\epsilon\hat{\phi}^{\star}(x_{1}^{i})$,
3   where $\hat{\phi}^{\star}(x_{1})$ from (1) given $x_{0}^{\text{true}}$
4 end for
5for $i\leftarrow 1$ to $N_{s}$ do
6 $\varphi_{1}(i)=\log p(z_{1},x_{1}^{i}|x_{0}^{\text{true}})$
7 end for
// 2) Sequential Particle Update
8 for $t=2,\dots,T$ do
9 for $i=1,2,\dots,N_{s}$ in parallel do
10    $x_{t}^{i}\leftarrow x_{t}^{i}+\epsilon\hat{\phi}^{\star}(x_{t}^{i})$,
11      where $\hat{\phi}^{\star}(x_{t})$ from (IV.3) given $\{x_{t-1}^{j}\}_{j=1}^{N_{s}}$
12   end for
 // 3) Forward Recursion
13 for $i\leftarrow 1$ to $N_{s}$ do
[…]
16   end for
17
18 end for
// 4) Backtracking
19 $i_{T}=\arg\!\max_{i}\varphi_{T}(i)$
20 $\hat{x}^{\text{MAP}}_{T}=x^{i_{T}}_{T}$
21 for $t=T-1,\dots,1$ do
22 $i_{t}=\psi_{t+1}(i_{t+1})$
23 $\hat{x}^{\text{MAP}}_{t}=x_{t}^{i_{t}}$
24 end for
[…]
Algorithm 2 Stein-MAP-Seq
In summary, Stein-MAP-Seq achieves efficient MAP sequence estimation through the synergistic combination of SVGD’s flexible particle-based approximation and dynamic programming’s optimal substructure exploitation. The method transforms the intractable continuous optimization into a manageable discrete problem while preserving the temporal dependencies essential for accurate sequential estimation.
@@ V Demonstrations
In this section, we rigorously evaluated the proposed method on challenging multimodal scenarios, including (A) nonlinear dynamics with ambiguous measurements, (B) pose estimation under unknown data associations, and (C) real-world range-only (wireless) localization under temporary unobservability.
@@
  • Extended Kalman Filter (EKF) [39] and Particle Filter (PF) [40], both MMSE estimators, and Extended Kalman Smoother (EKS), a MMSE sequence estimator;
  • Iterated EKF (iEKF), a (local) MAP estimator [^4] The iterative procedure can be interpreted as the Gauss-Newton method for finding the MAP estimate at the correction step [41]., and Iterated EKS (iEKS), a (local) MAP sequence estimator;
@@
All methods were implemented in Python and run on a laptop with an Intel Core i7‑10510U CPU and 32GB of RAM. For the resampling step of the PF, we applied stratified sampling at every step. For the Stein series, we set $\epsilon=0.001$ and fixed the number of iterations to $1000$ for all scenarios.
@@
with $v_{t-1}\sim\mathcal{N}(0,5)$, $r_{t}\sim\mathcal{N}(0,10)$ and $x_{0}\sim\mathcal{N}(0,5)$. The significance of this example is that the term $x_{t}^{2}$ in (22b) introduces a sign ambiguity, e.g., $x_{t}=+a$ and $x_{t}=-a$ yield the same measurement, resulting in a bimodal posterior (see Fig. 2, top)
@@
To evaluate the state estimation methods, we generated $50$ independent trajectories of $100$ time steps each. For each estimator, we computed the root‑mean‑square error (RMSE) of the state estimates, averaged over all Monte Carlo runs. The results in Fig. 2, bottom, Fig. 3, and Table I demonstrated that Stein-MAP-Seq significantly outperformed several state estimation methods, indicating that Stein-MAP-Seq is robust to bimodal posterior distributions. This performance advantage can be attributed to the mode-seeking behavior of SVGD, even with a small number of particles, and the use of Viterbi-style dynamic programming to identify the globally optimal trajectory over the sampled space.
Figure 2: (Top) True state and noisy observations (red dots) with sign ambiguity, resulting in a bimodal posterior distribution. (Bottom) One Monte Carlo simulation result. The simulation showed that Stein-MAP-Seq (cyan solid line with markers) significantly outperformed several state estimation methods, demonstrating robustness to bimodal posterior distributions.
Figure 3: Results from $50$ Monte Carlo simulations over $100$ steps under nonlinear dynamics with ambiguous measurements. In sampling-based methods, the number following each method name indicates the particle count; in Gaussian-based methods, it indicates the number of iterations.
@@
The methods, such as EKF, EKS, iEKF-3, and iEKS-3, rely on Gaussian (unimodal) assumptions and local linearization in filtering and smoothing, making them prone to converge to locally optimal trajectories. In this setup, iEKF-3 and iEKS-3 were initialized with EKF and iEKF-3 estimates, respectively, and refined the trajectory through iterative forward–backward local linearization. The number following the method name denotes the iteration number. iEKF-3 and iEKS-3 showed worse performance than EKF and EKS because their initial trajectories were suboptimal, and in the presence of a bimodal posterior, local optimization reinforced incorrect estimates.
For PF-MAP-Seq, even with a large number of particles ($100$ or $500$) relying on random sampling, its mode-seeking capability was weaker than that of Stein-MAP-Seq, which used significantly fewer particles ($10$ or $50$). In the case of PF-MAP and SPF-MAP, the estimators converged to an incorrect mode due to sign ambiguity in the observations and the lack of temporal dependency modeling, operating on points rather than entire trajectories. This led to the worst estimation performance, even worse than MMSE estimators such as EKF, PF, and SPF.
TABLE I: Robustness Evaluation for Scenario (A), Averaged RMSE
EKF EKS $N_{s}$ PF PF-MAP PF-MAP-Seq
1e2 $6.0441$ $8.1799$ $6.7130$
$6.7678$ $\mathbf{6.0377}$ 5e2 $6.0071$ $8.2770$ $6.8770$
1e3 $\mathbf{5.9816}$ $8.2658$ $6.9237$
iEKF-3 iEKS-3 $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
1e1 $4.4294$ $8.6393$ $\mathbf{2.1058}$
$7.3307$ $7.6991$ 5e1 $4.3164$ $8.6605$ $2.1881$
1e2 $4.3127$ $8.6848$ $2.3284$
Among the Stein variants, SPF is an MMSE estimator, while SPF-MAP provides a point-wise MAP estimate. Both are outperformed by Stein-MAP-Seq, which explicitly models temporal dependencies across the trajectory. Interestingly, for Stein-MAP-Seq, using $10$ particles results in better performance than using $50$ or $100$ particles. In SVGD, each particle update is influenced by all other particles through the kernel term (the second term in (IV.3)), and excessive repulsion can push particles into low-density regions to maintain diversity, which can negatively affect mode-seeking behavior.
@@ V-B Pose Estimation under Unknown Data Association
In the second example, we consider pose estimation on SE(2) under unknown data association. The discrete‐time pose is updated by
@@
with the pose process noise $\mathcal{N}(\mathbf{0},\mathbf{Q}_{t-1})$ where $\mathbf{Q}_{t-1}=\mathrm{diag}\bigl{[}(\alpha_{x}\,\Delta t)^{2},\;(\alpha_{y}\,\Delta t)^{2},\;(\alpha_{\theta}\,\Delta t)^{2}\bigr{]}$. The state vector is $\mathbf{x}_{t}=\bigl{[}x_{t},\;y_{t},\;\theta_{t}\bigr{]}$ (m,m,rad)(m,\,m,\,rad) with control inputs $\bigl{[}v_{t},\;\omega_{t}\bigr{]}$ (m/s,rad/s)(m/s,\,rad/s), and $\Delta t$ denotes the sampling period. Here, $\alpha_{x}$ and $\alpha_{y}$ quantify translational noise per unit time along the body‑fixed $x$ and $y$ axes, and $\alpha_{\theta}$ quantifies rotational noise per unit time. In our simulation, we use $\alpha_{x}=\alpha_{y}=0.05$, $\alpha_{\theta}=0.035$, $\Delta t=0.1$, and initialize the covariance as $\Sigma_{0}=\mathrm{diag}[0.2^{2},\,0.2^{2},\,0.035^{2}]$.
@@
where $z^{r}_{t,l}\in\mathbb{R}_{\geq 0}\,(m)$, $z^{b}_{t,l}\in[-\pi,\pi]\,(rad)$, $\Gamma_{l}=[\Gamma_{l,x},\Gamma_{l,y}]^{\top}\in{\mathbb{R}}^{2}$ is the $l$-th anchor position, and $P_{t}=[x_{t},\,y_{t}]^{\top}\in{\mathbb{R}}^{2}$ is the robot’s position at time $t$. Here, $\|\cdot\|_{2}$ denotes the Euclidean norm, and the measurement noises are $r^{r}_{t,l}\sim\mathcal{N}(0,1^{2})$ and $r^{b}_{t,l}\sim\mathcal{N}(0,0.17^{2})$.
@@
$p(\mathbf{z}_{t}\mid\mathbf{x}_{t})=\sum_{l=1}^{4}p\bigl{(}\mathbf{z}_{t}\mid\mathbf{x}_{t},l\bigr{)}\,p(l)\,$, (25)
@@
where $p(l)=\tfrac{1}{4}$. This results in an equally weighted mixture of the four single‑landmark likelihoods, producing a multimodal posterior distribution with multiple peaks (similar to a Gaussian mixture model).
Figure 4: One Monte Carlo simulation: a robot receives observations from 4 landmarks with unknown data associations, producing a multi-peak posterior. Stein-based estimation outperformed Gaussian-based and PF-based estimation.
Figure 5: Results from $10$ Monte Carlo simulations over $630$ steps for pose estimation under unknown data association. Methods are ordered from left to right by increasing RMSE.

### v4

@@ IV-A Sequential Variational Inference for Dynamical Systems
Under Assumption IV.1, the following lemma establishes the key structural result that distinguishes our approach from standard VI methods.
@@
Significance of Lemma IV.1: This result is not a standard VI factorization—it establishes a sequential conditional structure that fundamentally differs from both parametric VI and mean field approximations. Unlike mean field approaches that sacrifice temporal structure through independence ($q(x_{0:T})=\prod_{t}q(x_{t})$), the factorization in (13) preserves temporal dependencies through conditional distributions $q(x_{t}|x_{t-1})$ while enabling sequential recursive computation. This structure is essential for: (i) justifying the SVGD particle updates in Eq. (IV.3) that condition on particles from previous time steps, (ii) deriving the factorized ELBO in Lemma IV.2 with marginalized conditional KL terms, and (iii) enabling Viterbi-style dynamic programming over time-evolving particle distributions.
According to Lemma IV.1, the proposal distribution $q(x_{0:T})$ in (IV-A) admits this factorized structure. Consequently, the ELBO can be expressed as a summation of factorized KL divergence terms, each marginalized with respect to the previous optimal distribution, as stated in the following result.
@@ IV-B1 Viterbi-Style Dynamic Programming Algorithm
IV-B2 Theoretical Guarantees
@@
1
1ex// 1) Initialization $(t=1)$
2 for $i=1,2,\dots,N_{s}$ in parallel do
3 $x_{1}^{i}\leftarrow x_{1}^{i}+\epsilon\hat{\phi}^{\star}(x_{1}^{i})$,
4   where $\hat{\phi}^{\star}(x_{1})$ from (1) given $x_{0}^{\text{true}}$
5 end for
6for $i\leftarrow 1$ to $N_{s}$ do
7 $\varphi_{1}(i)=\log p(z_{1},x_{1}^{i}|x_{0}^{\text{true}})$
8 end for
// 2) Sequential Particle Update
9 for $t=2,\dots,T$ do
10 for $i=1,2,\dots,N_{s}$ in parallel do
11    $x_{t}^{i}\leftarrow x_{t}^{i}+\epsilon\hat{\phi}^{\star}(x_{t}^{i})$,
12      where $\hat{\phi}^{\star}(x_{t})$ from (IV.3) given $\{x_{t-1}^{j}\}_{j=1}^{N_{s}}$
13   end for
 // 3) Forward Recursion
14 for $i\leftarrow 1$ to $N_{s}$ do
[…]
17   end for
18
19 end for
// 4) Backtracking
20 $i_{T}=\arg\!\max_{i}\varphi_{T}(i)$
21 $\hat{x}^{\text{MAP}}_{T}=x^{i_{T}}_{T}$
22 for $t=T-1,\dots,1$ do
23 $i_{t}=\psi_{t+1}(i_{t+1})$
24 $\hat{x}^{\text{MAP}}_{t}=x_{t}^{i_{t}}$
25 end for
[…]
Algorithm 2 Stein-MAP-Seq
@@ 24 $\hat{x}^{\text{MAP}}_{t}=x_{t}^{i_{t}}$
Theorem IV.1, combined with Lemma IV.3, establishes rigorous convergence guarantees for Stein-MAP-Seq. Lemma IV.3 ensures that the sequential SVGD updates provably decrease the marginalized conditional KL divergence at each time step, guaranteeing that the generated particles $\{x_{t}^{i}\}_{i=1}^{N_{s}}$ concentrate around the modes of the target conditional posteriors $p(z_{t},x_{t}|x_{t-1})$. Theorem IV.1 then establishes that the Viterbi-style dynamic programming over these particle sets recovers the trajectory that maximizes the ELBO, which converges to the true log-posterior $J(s)$ as $\varepsilon\to 0$. Together, these results provide end-to-end theoretical justification: the SVGD-generated particles provably approximate optimal conditional distributions (Lemma IV.3), and the dynamic programming algorithm provably identifies the globally optimal MAP trajectory within this discrete particle support (Theorem IV.1), with approximation error $|\mathcal{L}_{\varepsilon}(s)-J(s)|\to 0$ as particle quality improves.
Having established the theoretical foundations and convergence guarantees, we now analyze the computational complexity of Stein-MAP-Seq.
IV-C Computational Analysis of Stein-MAP-Seq
Table I summarizes the computational complexity of the main components of Stein-MAP-Seq. The dominant cost in the SVGD update arises from the kernel interaction term, which scales quadratically with the number of particles $N_{s}$ and linearly with the state dimension $n_{x}$, resulting in an $\mathcal{O}(N_{s}^{2}n_{x})$ complexity. Although this double summation in Eq. (IV.3) can be computationally demanding, it is consistent with standard SVGD-based methods and, importantly, is fully parallelizable. Both the log-posterior gradient evaluations and the pairwise kernel interactions are independent across particles, making them well-suited for efficient GPU acceleration. To complement this theoretical analysis, Appendix -F reports GPU-based per-step runtime evaluations of (IV.3) as a function of the number of particles and state dimension (Fig. 14), demonstrating stable and predictable runtime behavior in practice.
Beyond parallelization, Stein-MAP-Seq further mitigates computational cost through the mode-seeking yet dispersive behavior of SVGD. By concentrating particles around multiple dominant modes while maintaining separation between them, SVGD produces a compact set of high-quality candidates for MAP trajectory estimation. As a result, Stein-MAP-Seq operates reliably with a significantly smaller number of particles than MC sampling–based MAP-Seq methods.
The primary computational bottleneck in MAP sequence estimation arises from the forward recursion in (21), which incurs an $\mathcal{O}(N_{s}^{2})$ cost per time step. Crucially, the reduction in the effective particle count $N_{s}$ directly translates into a substantial reduction in this dominant cost. Moreover, the forward recursion remains amenable to parallelization through independent score computations followed by max-reduction operations, and its complexity is independent of the state dimension $n_{x}$. The backtracking step scales linearly with the trajectory length $T$ and contributes negligible overhead. Consequently, Stein-MAP-Seq achieves computationally efficient MAP sequence estimation by combining parallelizable SVGD updates with particle-efficient mode discovery, effectively reducing the cost of the dominant forward recursion even in high-dimensional systems. This distinction explains why the practical efficiency of Stein-MAP-Seq stems not from asymptotic complexity reduction, but from particle efficiency and parallel execution.
Overall, the above analysis highlights that the computational efficiency of Stein-MAP-Seq is closely tied to its parallel-friendly structure. As a result, Stein-MAP-Seq is naturally suited for efficient execution on parallel computing architectures. Although SVGD requires synchronization to aggregate kernel interactions, this synchronization is limited to standard reduction operations and does not involve resampling or particle selection. In contrast, Monte Carlo sampling methods such as particle filters [41] rely on resampling procedures that introduce non-deterministic state selection, resulting in variable computation paths across runs. Moreover, particles cannot be processed independently because resampling depends on global statistics, such as normalized weights and cumulative sums, which significantly limit parallel efficiency in practice.
TABLE I: Computational complexity of Stein-MAP-Seq components
Component Time Complexity Parallelizable [^4] The parallelizable column indicates execution structure and does not change the asymptotic computational complexity.
Log-posterior gradient (in Eq. (IV.3)) $\mathcal{O}(N_{s}n_{x})$ Yes (particle-wise independent)
Kernel interaction (in Eq. (IV.3)) $\mathcal{O}(N_{s}^{2}n_{x})$ Yes (pairwise independent; reduction)
Forward recursion (Eq. (21)) $\mathcal{O}(N_{s}^{2})$ Yes (per time step; max-reduction)
Backtracking $\mathcal{O}(T)$ No (time-serial)
@@ V Demonstrations
In this section, we rigorously evaluated the proposed method on challenging multimodal scenarios, including (A) nonlinear dynamics with ambiguous measurements, (B) pose estimation under unknown data associations with outliers, (C) range-only (wireless) localization under temporary unobservability, and (D) high-dimensional manipulator with kinematic redundancy.
@@
  • Extended Kalman Filter (EKF) [42] and Particle Filter (PF) [43], both MMSE estimators, and Extended Kalman Smoother (EKS), a smoothing-based MMSE sequence estimator;
  • Iterated EKF (iEKF), a local MAP estimator [^5] The iterative procedure can be interpreted as the Gauss-Newton method for finding the MAP estimate at the correction step [44]., and Iterated EKS (iEKS), a local MAP sequence estimator;
  • GTSAM [45], batch MAP sequence estimator via factor-graph optimization;
@@
All methods were implemented in Python and evaluated on a laptop equipped with an Intel Core i7-10510U CPU and 32 GB of RAM. Our method was implemented in PyTorch and executed on a GPU to accelerate the parallel particle updates in (IV.3). This GPU acceleration affects only computational speed and does not modify the algorithmic structure or estimation results. For PF-based methods, stratified resampling was applied at each time step. For the Stein-based methods, the step size was set to $\epsilon=0.005$, and the number of iterations was fixed to $100$ across all scenarios. We employ a radial basis function (RBF) kernel to model smooth interactions among particles. For GTSAM, we adopt the Levenberg–Marquardt optimizer with a maximum of 50 iterations to improve robustness under strong nonlinearities.
[…]
@@
with $v_{t-1}\sim\mathcal{N}(0,5)$, $r_{t}\sim\mathcal{N}(0,4^{2})$, and $x_{0}\sim\mathcal{N}(0,5)$. The significance of this example is that the term $x_{t}^{2}$ in (22b) introduces a sign ambiguity, e.g., $x_{t}=+a$ and $x_{t}=-a$ yield the same measurement, resulting in a bimodal posterior (see Fig. 2, middle). Moreover, the large process uncertainty spreads the predicted prior across competing modes, while the high measurement uncertainty weakens observational correction, preventing reliable local mode discrimination.
@@
To evaluate the state estimation methods, we generated $50$ independent trajectories of $100$ time steps each. For each estimator, we computed the root‑mean‑square error (RMSE) of the state estimates, averaged over all Monte Carlo runs. The results in Fig. 2, bottom, Fig. 3, and Table II demonstrate that Stein-MAP-Seq consistently outperformed competing estimators, highlighting its robustness to bimodal posterior distributions under high process and measurement uncertainty. This performance advantage arises from the complementary interaction between SVGD’s mode-seeking behavior and Viterbi-style dynamic programming. SVGD concentrates particles around dominant posterior modes using only a small particle set, while Viterbi’s forward recursion exploits this focused representation to recover the globally optimal trajectory over the sampled space.
Figure 3: RMSE results over $50$ Monte Carlo simulations with $100$ time steps are reported for a bimodal posterior under high process and measurement uncertainty. Finite-state estimators are generally more robust than Gaussian-assumed estimators in multimodal settings, except for point-wise MAP estimators. Point-wise MAP estimators frequently select incorrect modes and therefore perform worse than MMSE estimators, which effectively average across modes. Among MAP sequence estimators, Stein-MAP-Seq and GTSAM achieve the best performance, whereas PF-based MAP methods exhibit the highest RMSE due to unstable mode selection.
@@
The methods, such as EKF, EKS, iEKF-3, and iEKS-3, rely on Gaussian (unimodal) assumptions and local linearization, making them prone to converge to locally optimal trajectories. In this setup, large process noise spreads the prior across competing modes, while high measurement uncertainty weakens observational correction, preventing both filtering and smoothing from resolving the sign ambiguity. iEKF-3 and iEKS-3 were initialized with EKF and iEKF-3 estimates, respectively, and refined the trajectory through iterative forward–backward local linearization, where the number following the method name denotes the iteration count. However, because the initial trajectories were already biased toward incorrect modes, the subsequent local optimization in iEKF-3 and iEKS-3 reinforced these errors, resulting in worse performance than EKF and EKS, as shown in Table II.
TABLE II: Robustness Evaluation for Scenario (A), Averaged RMSE
Gaussian-assumed Estimation Finite-state Estimation
EKF EKS GTSAM $N_{s}$ PF PF-MAP PF-MAP-Seq
5e2 $6.1288$ $8.1654$ $7.0949$
$7.2289$ $6.3422$ $2.5564$ 1e3 $6.1320$ $8.1829$ $7.0696$
2e3 $\mathbf{6.1281}$ $8.1863$ $7.1363$
iEKF-3 iEKS-3 GTSAM-Stein $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
1e1 $4.3719$ $8.8307$ $\mathbf{2.0462}$
$7.9307$ $8.1320$ $\mathbf{1.5509}$ 2e1 $4.4282$ $8.6583$ $2.1447$
4e1 $4.4964$ $8.8143$ $2.2761$
GTSAM performed batch optimization over the entire trajectory. However, under high process and measurement uncertainty, the resulting factor graph became weakly constrained, making the optimization highly sensitive to initialization. As a result, GTSAM frequently converged to an incorrect mode (see Fig. 2, bottom), yielding suboptimal trajectory estimates. When initialized with the trajectory estimate obtained from Stein-MAP-Seq (see Table II), GTSAM-Stein consistently converged to the correct mode, demonstrating the effectiveness of Stein-MAP-Seq as an initialization strategy for nonconvex batch MAP optimization.
For PF-MAP-Seq, even with a large number of particles ($\geq 500$) drawn via random sampling, its mode-seeking capability was markedly weaker than that of Stein-MAP-Seq, which operated with significantly fewer particles ($\leq 50$). This behavior arose from particle path degeneracy induced by resampling and sampling variability under a bimodal posterior, where importance sampling failed to concentrate particles around a single dominant trajectory consistently. As a result, MAP decoding became unstable and prone to mode switching, producing zig-zag trajectories (see Fig. 2, bottom). In contrast, SVGD promoted coherent particle concentration around dominant modes and enabled reliable MAP sequence recovery. For PF-MAP and SPF-MAP, the estimators converged to incorrect modes due to the absence of explicit temporal dependency modeling, as they operated on marginal state estimates rather than entire trajectories. Consequently, their estimation performance was the worst among all methods, even inferior to MMSE-based estimators such as EKF, PF, and SPF.
Figure 4: The averaged computation time per step is reported over $50$ Monte Carlo simulations. Filtering-based methods are more computationally efficient than MAP-Seq estimators, since MAP-Seq estimation requires trajectory-level inference. PF-MAP-Seq estimators incur the highest computational cost due to the large number of particles required, whereas Stein-MAP-Seq achieves competitive performance with fewer particles and computational cost comparable to GTSAM.
Among the Stein variants, Stein-MAP-Seq exhibited superior performance compared to SPF and SPF-MAP, despite their explicit modeling of temporal dependencies across the trajectory. Notably, Stein-MAP-Seq achieved higher estimation accuracy using significantly fewer particles. This advantage arises because, unlike MMSE-oriented estimators that require accurate approximation of the full posterior distribution, MAP-Seq estimation benefited from the mode-seeking behavior and the repulsive interactions in (IV.3), enabling reliable forward recursion and, consequently, accurate MAP trajectory recovery.
As shown in Fig. 4, filtering-based methods incurred the lowest computation time, whereas MAP-Seq estimators were inherently more expensive due to trajectory hypothesis management. Among finite-state estimators, the Stein series maintained a moderate computational cost with online forward recursion that scaled favorably with the number of particles, remaining comparable to GTSAM in practice while significantly outperforming PF-based MAP estimators. In contrast, PF-MAP-Seq exhibited rapidly increasing computation time as the particle count grew, reflecting the high computational burden of forward recursion over a large set of MAP candidate trajectories. Overall, these results indicated that Stein-MAP-Seq achieved a favorable accuracy–computation trade-off with a reduced number of particles required for reliable MAP-Seq estimation.
@@ V-B Pose Estimation under Unknown Data Association
In the second example, we consider pose estimation on SE(2) under unknown data association with outliers. The discrete‐time pose is updated by
@@
with the pose process noise $\mathcal{N}(\mathbf{0},\mathbf{Q}_{t-1})$ where $\mathbf{Q}_{t-1}=\mathrm{diag}\bigl[(\alpha_{x}\,\Delta t)^{2},\;(\alpha_{y}\,\Delta t)^{2},\;(\alpha_{\theta}\,\Delta t)^{2}\bigr]$. The state vector is $\mathbf{x}_{t}=\bigl[x_{t},\;y_{t},\;\theta_{t}\bigr]^{\top}$ (m,m,rad)(m,\,m,\,rad) with control inputs $\bigl[v_{t},\;\omega_{t}\bigr]^{\top}$ (m/s,rad/s)(m/s,\,rad/s), and $\Delta t$ denotes the sampling period. Here, $\alpha_{x}$ and $\alpha_{y}$ quantify translational noise per unit time along the body‑fixed $x$ and $y$ axes, and $\alpha_{\theta}$ quantifies rotational noise per unit time. In our simulation, we use $\alpha_{x}=\alpha_{y}=0.3$, $\alpha_{\theta}=0.105$, $\Delta t=0.1$, and initialize the covariance as $\Sigma_{0}=\mathrm{diag}(0.2^{2},\,0.2^{2},\,0.035^{2})$.
@@
where $z^{r}_{t,l}\in\mathbb{R}_{\geq 0}\,(m)$, $z^{b}_{t,l}\in[-\pi,\pi]\,(rad)$, $\Gamma_{l}=[\Gamma_{l,x},\Gamma_{l,y}]^{\top}\in{\mathbb{R}}^{2}$ is the $l$-th anchor position, and $P_{t}=[x_{t},\,y_{t}]^{\top}\in{\mathbb{R}}^{2}$ is the robot’s position at time $t$. Here, $\|\cdot\|_{2}$ denotes the Euclidean norm, and the measurement noises are $r^{r}_{t,l}\sim\mathcal{N}(0,1^{2})$ and $r^{b}_{t,l}\sim\mathcal{N}(0,0.44^{2})$.
@@
$p(\mathbf{z}_{t}\mid\mathbf{x}_{t})=\sum_{l=1}^{4}\nolimits p\bigl(\mathbf{z}_{t}\mid\mathbf{x}_{t},l\bigr)\,p(l)\,$, (25)
@@
where $p(l)=1/4$. This formulation results in an equally weighted mixture of the four single-landmark likelihoods, producing a multimodal posterior distribution with multiple peaks, analogous to a Gaussian mixture model.
(a) Multimodal Likelihoods
(b) Gaussian-assumed Estimation
(c) PF-based Estimation
(d) Stein-based Estimation
Figure 5: (a) A robot follows a counter-clockwise reference trajectory (black dashed curve) while executing noisy motion dynamics at every time step, causing the realized trajectory to gradually deviate from the noise-free reference. At each step, the robot receives a range–bearing measurement from one of four landmarks with unknown data association and heavy-tailed outliers, resulting in a highly multimodal measurement likelihood that varies across locations. (b) Trajectory estimates produced by Gaussian-assumed methods, illustrating bias and inconsistency under nonlinear motion and multimodal measurement likelihoods. (c) PF-based estimators with 500 particles improve robustness compared to Gaussian methods but still exhibit noticeable trajectory distortion due to weight degeneracy and resampling effects. (d) Stein-based estimators with 20 particles accurately recover the reference trajectory while remaining robust to multimodality and outliers, demonstrating reliable estimation with a small number of particles.
Figure 6: Results from $10$ Monte Carlo simulations over $630$ time steps are reported for pose estimation under unknown data association with heavy-tailed outliers. Stein-based methods demonstrate improved robustness and lower estimation error, whereas PF-based and Gaussian methods show increased sensitivity to ambiguous data association and heavy-tailed noise. Among Gaussian-assumed estimators, GTSAM benefits from trajectory-level smoothing, leading to improved performance relative to filtering-based approaches despite its reliance on unimodal assumptions.
To evaluate robustness under multimodal uncertainty, we consider a long-horizon localization task with nonlinear motion dynamics and constant turning. Translational and rotational process noise are injected at every time step, resulting in gradual uncertainty accumulation over time. At each step, the observed landmark is randomly selected, inducing abrupt changes in the measurement model and significantly reducing the temporal consistency of observations. In addition, measurement noise follows a heavy-tailed mixture distribution, introducing large outliers ($\times 4$) with a $10\%$ probability in both range and bearing measurements. Collectively, these factors yield a likelihood function that is highly multimodal and heavy-tailed over time, posing substantial challenges for accurate and robust state estimation.